# MIMO Detection with Spatial Sigma-Delta ADCs: A Variational Bayesian Approach

### v1

@@ II-B1 Few-Bit ADCs with Conventional Quantizers
where $\mathcal{Q}_{b}(.)$ denotes the $b$-bit ADC operation applied element-wise to its vector argument. The uniform quantizer $\mathcal{Q}_{b}(.)$ performs $b$-bit scalar quantization which produces $12^{b}-$ 1 quantized output levels belonging to the set $\{d_{1},\ldots,d_{2^{b}-1}\}$. Without loss of generality, we assume ${-\infty=d_{0}<d_{1}\ldots<d_{2^{b}-1}<d_{2^{b}}=+\infty}$. The quantization thresholds are given by [10]
$d_{m}=\big{(}-2^{b-1}+m\big{)}\Lambda,\quad\forall m\in\mathcal{M}={1,\ldots,2^{b-1}},$ (6)
@@ II-B2 Few-Bit ADCs with Sigma-Delta Quantizers
where $x_{i}\triangleq{\bf h}_{i}^{T}{\bf s}+n_{i}$ is the unquantized received signal at antenna $i$, $r_{0}$ and $y_{0}$ are set to 00, and $\phi$ denotes a phase shift applied to the signal between two adjacent antennas.
Figure 1: Illustration of the spatial $1^{\rm st}$-order $\Sigma\Delta$ architecture at an $N$-antenna receiver.
@@
In the following, we present a linearized model for spatial $1^{\rm st}$-order $\Sigma\Delta$ ADCs and the corresponding linear minimum mean squared error (LMMSE) receiver to estimate $\mathbf{s}$. Subsequent developments will include the design of efficient VB-based MIMO detection algorithms for $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ ADCs.
III Linear Detection with Spatial $\Sigma\Delta$ Quantized Observations
Using the Bussgang decomposition [38], the output of the $\Sigma\Delta$ array in (9) in vector form is linearized as [20, Eq. (16)]:
$\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{U}^{-1}\mathbf{q},$ (10)
where $\mathbf{q}=[q_{1},\ldots,q_{N}]^{T}$ represents the effective quantization noise, and $\mathbf{U}$ is defined as
For the special case of one-bit $\Sigma\Delta$ quantization, the variance of $q_{i}$ is given by [20, Eq. (33)]
@@ III Linear Detection with Spatial $\Sigma\Delta$ Quantized Observations
Based on the linearized model in (10), an LMMSE receiver $\mathbf{W}_{\rm LMMSE}$ that minimizes the mean squared error ${\mathbb{E}\big{[}\|\mathbf{s}-\mathbf{Wy}\|^{2}\big{]}}$ in estimating $\mathbf{s}$ can be expressed as
$\mathbf{W}_{\rm LMMSE}=\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}(\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}+N_{0}\mathbf{I}_{N}+\mathbf{U}^{-1}\bm{\Sigma}_{\mathbf{q}}\mathbf{U}^{-H})^{-1},$ (13)
where $\bm{\Sigma}_{\mathbf{s}}$ and $\bm{\Sigma}_{\mathbf{q}}=\textrm{diag}(\tau_{q_{1}},\ldots,\tau_{q_{N}})$ are the covariance matrices of $\mathbf{s}$ and $\mathbf{q}$, respectively. As mentioned earlier, the accuracy of the Bussgang decomposition depends on the Gaussianity of the input to the quantization non-linearity, which does not always hold, particularly when the input symbol is discrete. To overcome this challenge, in the next section, we introduce an alternative MIMO detection approach based on VB inference to handle the nonlinear $\Sigma\Delta$ receiver architecture.
IV Variational Bayes for Few-Bit $1^{\rm st}$-Order $\Sigma\Delta$ MIMO Detection
@@
IV-A Background on VB Inference
We denote the set of all observed variables as $\mathbf{y}$ and the set of $V$ latent variables as $\mathbf{x}$. To find a Bayes estimate of $\mathbf{x}$, one would need to determine the posterior distribution $p(\mathbf{x}|\mathbf{y})$, which is often computationally intractable. To overcome this challenge, the VB method aims to find a distribution $q(\mathbf{x})$ characterized by variational parameters within a predefined family $\mathcal{Q}$ of densities, such that $q(\mathbf{x})$ closely approximates $p(\mathbf{x}|\mathbf{y})$. To this end, VB defines an optimization problem leveraging the Kullback-Leibler (KL) divergence from $q(\mathbf{x})$ to $p(\mathbf{x}|\mathbf{y})$:
$q^{\star}(\mathbf{x})=$ $\arg\min_{q(\mathbf{x})\in\mathcal{Q}}\text{KL}\big{(}q(\mathbf{x})\|p(\mathbf{x}|\mathbf{y})\big{)},$ (14)
$=$ $\arg\min_{q(\mathbf{x})\in\mathcal{Q}}\big{\{}\mathbb{E}_{q(\mathbf{x})}\big{[}\ln q(\mathbf{x})\big{]}-\mathbb{E}_{q(\mathbf{x})}\big{[}\ln p(\mathbf{x}|\mathbf{y})\big{]}\big{\}},$
where $q^{\star}(\mathbf{x})$ denotes the optimal variational distribution. The VB method assumes that $q(\mathbf{x})$ is drawn from the mean-field variational family [39], such that
$q(\mathbf{x})=\prod_{i=1}^{V}q_{i}(x_{i}),$ (15)
where the latent variables are mutually independent, with each being governed by a distinct factor in the variational density. The optimal value of $q_{i}(x_{i})$ is obtained as [39, Chapter 10]
$q_{i}^{\star}(x_{i})$ $\propto\mathrm{exp}\left\{\big{\langle}{\ln p(\mathbf{y},\mathbf{x})\big{\rangle}}\right\},$ (16)
[…]
IV-B VB Inference for Few-bit $\Sigma\Delta$ MIMO Detection
@@
This integral is intractable to evaluate in closed-form due to the high dimensional integration and the intricate correlation between $\mathbf{r}$ and $\mathbf{s}$. This challenge motivates us to explore a novel VB inference method to solve the data detection problem with $\Sigma\Delta$ arrays.
@@
We consider the residual interference-plus-noise as an unknown parameter $N_{0}^{\rm post}$, which must be estimated using the VB method. The dependency between random variables under spatial $\Sigma\Delta$ processing can be graphically modeled through the Bayesian network in Fig. 2, where $\gamma\triangleq 1/N_{0}^{\rm post}$ denotes the precision that is floated as an unknown random variable and the arrows indicate the conditional probability between variables. The main goal is to infer the distribution of data $\mathbf{s}$ given the observation $\mathbf{y}$. To accomplish this, we employ the mean-field variational distribution $q(\mathbf{s},\mathbf{r},\gamma)$, such that
@@
×$\;\;\times\prod_{i=1}^{N}p(y_{i}|r_{i})p(r_{i}|r_{i-1},y_{i-1},\mathbf{s},\gamma;\mathbf{H}),$ (19)
@@
1) Updating $r_{i}$: For $i=1,\ldots,N-1$, the variational distribution $q(r_{i})$ can be obtained by taking the expectation of the conditional in (IV-B) w.r.t. $q(\mathbf{s},\gamma)$ as
@@
Defining $v_{i}\triangleq(a_{i}+b_{i})/2$, we see that the variational distribution $q(r_{i})$ in (IV-B) is a truncated complex Gaussian distribution obtained by pruning the distribution $\mathcal{CN}\big{(}v_{i},1/(2\langle\gamma\rangle)\big{)}$ onto the interval $[y_{i}^{\mathrm{low}},y_{i}^{\mathrm{up}}]$. In Appendix A we show that the mean and variance of the resulting distribution are given by ${\langle r_{i}\rangle=\mathsf{F}_{r}(v_{i},2\langle\gamma\rangle,y_{i}^{\mathrm{low}},y_{i}^{\mathrm{up}})}$ and ${\tau_{r_{i}}={\sf G}_{r}(v_{i},2\langle\gamma\rangle,y_{i}^{\mathrm{low}},y_{i}^{\mathrm{up}})}$, respectively.
For $i=N$, while the factor $p(r_{N+1}|r_{N},y_{N},\mathbf{s},\gamma;\mathbf{H})$ does not exist, the variational distribution $q(r_{N})$ can be computed similarly to (IV-B) and is given by
@@
2) Updating $s_{k}$: The variational distribution $q(s_{k})$ is obtained by taking the expectation of the conditional in (IV-B) w.r.t. $q(\mathbf{r},\gamma)$:
@@
$\mathbf{z}_{k}=\mathbf{h}_{k}s_{k}+\mathcal{CN}(\mathbf{0},\langle\gamma\rangle^{-1}\mathbf{I}_{N})$ (26)
where $\mathbf{z}_{k}=[z_{1,k},\ldots,z_{N,k}]^{T}$. The variational mean and variance of $s_{k}$ can be expressed as $\langle s_{k}\rangle={\sf F}_{s}(\mathbf{z}_{k},\mathbf{h}_{k},\langle\gamma\rangle)$ and $\tau_{s_{k}}={\sf G}_{s}(\mathbf{z}_{k},\mathbf{h}_{k},\langle\gamma\rangle)$, respectively, as shown in Appendix B.
3) Updating $\gamma$: We assume a conjugate prior Gamma distribution $\rm Gamma(\alpha,\beta)$ for the precision $\gamma$, for given shape and rate parameters $\alpha$ and $\beta$, respectively. The variational distribution $q(\gamma)$ is obtained by taking the expectation of the conditional distribution in (IV-B) w.r.t. $q(\mathbf{s},\mathbf{r})$:
@@
where $u_{i}\triangleq\langle r_{i}\rangle-\mathbf{h}_{i}^{T}\langle\mathbf{s}\rangle-\mathrm{e}^{-j\phi}(\langle r_{i-1}\rangle-y_{i-1})$, $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}$ and $\bm{\Sigma}_{\mathbf{r}}={\rm diag}(\tau_{r_{1}},\ldots,\tau_{r_{N}})$ is the covariance matrix of $\mathbf{r}$. Here, $u_{i}$ denotes the residual term at receive antenna $i$, which would equals the noise term $n_{i}$ if $r_{i}$, $r_{i-1}$, and $\mathbf{s}$ were perfectly known.
The variational distribution $q(\gamma)$ is Gamma with mean
$\langle\gamma\rangle=\frac{N+\alpha}{\beta+\|\mathbf{u}\|^{2}+2\mathrm{Tr}\{\bm{\Sigma}_{\mathbf{r}}\}-\tau_{r_{N}}+\mathrm{Tr}\big{\{}\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}\big{\}}}.$ (29)
The proposed VB method for MIMO detection with $1^{\rm st}$-order $\Sigma\Delta$ quantization is summarized in Algorithm 1, where the parameter $\epsilon$ is the numerator of $\hat{\gamma}$ in the variational distributions of (IV-B) and (22) for updating $r_{i}$ and $r_{N}$, respectively.
Remark 1: In Algorithm 1, we use $\hat{r}_{i}^{t}$, $\hat{s}^{t}_{k}$, and $\hat{\gamma}^{t}$ to replace $\langle r_{i}\rangle$, $\langle s_{k}\rangle$, and $\langle\gamma\rangle$ at iteration $t$, and each iteration includes a round of updating the estimates of $\mathbf{r},\mathbf{s}$, and $\gamma$. Recalling $u_{i}=\langle r_{i}\rangle-\mathrm{e}^{-j\phi}(\langle r_{i-1}\rangle-y_{i-1})-\mathbf{h}_{i}^{T}\langle\mathbf{s}\rangle$, $a_{i}=\mathbf{h}_{i}^{T}\langle\mathbf{s}\rangle+\mathrm{e}^{-j\phi}(\langle r_{i-1}\rangle-y_{i-1})$ and $b_{i}=y_{i}+\mathrm{e}^{j\phi}\langle r_{i+1}\rangle-\mathrm{e}^{j\phi}\mathbf{h}_{i+1}^{T}\langle\mathbf{s}\rangle$, $a_{i}$ and $b_{i}$ can be rewritten as $a_{i}=\hat{r}_{i}^{t}-u_{i}$ and $b_{i}=\hat{r}_{i}^{t}+\mathrm{e}^{j\phi}u_{i+1}$ using the current estimates $\big{\{}\hat{r}_{i}^{t}\big{\}}$ and $\big{\{}\hat{\mathbf{s}}^{t}\big{\}}$. Thus, we have
@@
V Extension to MIMO Detection with $2^{\rm nd}$-Order $\Sigma\Delta$ Quantization

### v2

@@ II-B1 Few-Bit ADCs with Conventional Quantizers
where $\mathcal{Q}_{b}(.)$ denotes the $b$-bit ADC operation applied element-wise to its vector argument. The uniform quantizer $\mathcal{Q}_{b}(.)$ performs $b$-bit scalar quantization, which produces $2^{b}-1$ quantized output levels belonging to the set $\{d_{1},\ldots,d_{2^{b}-1}\}$. Without loss of generality, we assume $\infty=d_{0}<d_{1}\ldots<d_{2^{b}-1}<d_{2^{b}}=+\infty$}. The quantization thresholds are given by [10]
$d_{m}\!=\!\big(-2^{b-1}+m\big)\Lambda,\quad\forall m\in\mathcal{M}=\{1,\ldots,2^{b}-1\},$ (10)
@@ II-B2 Few-Bit ADCs with Sigma-Delta Quantizers
where $x_{i}\triangleq{\bf h}_{i}^{T}{\bf s}+n_{i}$ is the unquantized received signal at antenna $i$, $r_{0}$ and $y_{0}$ are set to 0, and $\phi$ denotes a phase shift applied to the signal between two adjacent antennas.
Figure 2: Illustration of the spatial $1^{\rm st}$-order $\Sigma\Delta$ architecture at an NN-antenna receiver.
@@
In the following, we present prior art on the linearized model for spatial $1^{\rm st}$-order $\Sigma\Delta$ ADCs and corresponding linear minimum mean-squared error (LMMSE) receiver to estimate $\mathbf{s}$. Subsequent developments will include the design of efficient VB-based MIMO detection algorithms for $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ ADCs.
II-C Linear Detection with Spatial $\Sigma\Delta$ Quantized Observations
Using the Bussgang decomposition [51], the output of the $\Sigma\Delta$ array in (13) in vector form is linearized as
${\mathbf{y}=\mathbf{x}+\mathbf{U}^{-1}\mathbf{q},}$ (14)
where $\mathbf{q}=[q_{1},\ldots,q_{N}]^{T}$ represents the effective quantization noise, and $\mathbf{U}$ is given in (59). A detailed derivation of (14) is provided in Appendix A. In (14), the unquantized signal $\mathbf{x}$ is separated from the shaped quantization noise $\mathbf{U}^{-1}\mathbf{q}$, which enables the use of linear detection methods for the $\Sigma\Delta$ array. For the special case of one-bit $\Sigma\Delta$ quantization, the variance of $q_{i}$ is given by [20, Eq. (33)]
@@ II-C Linear Detection with Spatial $\Sigma\Delta$ Quantized Observations
Based on the linearized model in (14), an LMMSE receiver $\mathbf{W}$ that minimizes the mean squared error $\mathbb{E}\big[\|\mathbf{s}-\mathbf{Wy}\|^{2}\big]$} in estimating $\mathbf{s}$ can be expressed as
$\mathbf{W}=\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}(\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}+N_{0}\mathbf{I}_{N}+\mathbf{U}^{-1}\bm{\Sigma}_{\mathbf{q}}\mathbf{U}^{-H})^{-1},$ (16)
where $\bm{\Sigma}_{\mathbf{s}}$ and $\bm{\Sigma}_{\mathbf{q}}=\textrm{diag}(\tau_{q_{1}},\ldots,\tau_{q_{N}})$ are the covariance matrices of $\mathbf{s}$ and $\mathbf{q}$, respectively.
If we take the mutual coupling effect into account, the LMMSE solution for estimating $\mathbf{s}$ can be obtained as
$\mathbf{W}^{\rm MC}=\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}(\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}+\mathbf{R}+\mathbf{U}^{-1}\bm{\Sigma}^{\rm MC}_{\mathbf{q}}\mathbf{U}^{-H})^{-1}\mathbf{y},$ (17)
where the covariance matrix of the quantization noise $\bm{\Sigma}^{\rm MC}_{\mathbf{q}}$ can be approximated as [52]
$\bm{\Sigma}^{\rm MC}_{\mathbf{q}}\simeq\mathrm{diag}(\mathbf{p}_{\mathbf{q}}),$ (18)
and
$\mathbf{p}_{\mathbf{q}}$ $=$ $\left(\frac{\pi}{2}\zeta-1\right)\bm{\Pi}\mathbf{p}_{\mathbf{x}},$ (19)
$\mathbf{p}_{\mathbf{x}}$ $=$ $\begin{bmatrix}\mathbb{E}\big[|x_{1}|^{2}\big],\mathbb{E}\big[|x_{2}|^{2}\big],\cdots,\mathbb{E}\big[|x_{N}|^{2}\big]\end{bmatrix}^{T},$ (20)
$\bf{\Pi}=$
$\begin{bmatrix}1&0&\ldots&&0&0\cr\left(\frac{\pi}{2}\zeta-1\right)&1&0&\ldots&&0\cr\vdots&\ddots&1&&&\vdots\cr\left(\frac{\pi}{2}\zeta-1\right)^{m}&\ddots&\ddots&\ddots&&\cr\vdots&\ddots&\ddots&\ddots&\ddots&\cr\left(\frac{\pi}{2}\zeta-1\right)^{M-1}&\cdots&\left(\frac{\pi}{2}\zeta-1\right)^{m}&\cdots&\left(\frac{\pi}{2}\zeta-1\right)&1\end{bmatrix}$, (21)
where $\zeta=1.13$ is a correction factor.
As mentioned earlier, the accuracy of the Bussgang decomposition depends on the Gaussianity of the input to the quantization non-linearity, which does not always hold, particularly when the input symbol is discrete. To overcome this challenge, in the next section, we introduce an alternative MIMO detection approach based on VB inference to handle the nonlinear $\Sigma\Delta$ receiver architecture.
III Variational Bayes for Few-Bit $1^{\rm st}$-Order $\Sigma\Delta$ MIMO Detection
@@
III-A Background on VB Inference
We denote the set of all observed variables as $\bar{\mathbf{y}}$ and the set of $V$ latent variables as $\bar{\mathbf{x}}$. To find a Bayes estimate of $\bar{\mathbf{x}}$, one would need to determine the posterior distribution $p(\bar{\mathbf{x}}|\bar{\mathbf{y}})$, which is often computationally intractable. To overcome this challenge, the VB method aims to find a distribution $q(\bar{\mathbf{x}})$ characterized by variational parameters within a predefined family $\mathcal{Q}$ of densities, such that $q(\bar{\mathbf{x}})$ closely approximates $p(\bar{\mathbf{x}}|\bar{\mathbf{y}})$. To this end, VB defines an optimization problem leveraging the Kullback-Leibler (KL) divergence from $q(\bar{\mathbf{x}})$ to $p(\bar{\mathbf{x}}|\bar{\mathbf{y}})$:
$q^{\star}(\bar{\mathbf{x}})=$ $\arg\min_{q(\bar{\mathbf{x}})\in\mathcal{Q}}\text{KL}\big(q(\bar{\mathbf{x}})\|p(\bar{\mathbf{x}}|\bar{\mathbf{y}})\big),$ (22)
$=$ $\arg\min_{q(\bar{\mathbf{x}})\in\mathcal{Q}}\big\{\mathbb{E}_{q(\bar{\mathbf{x}})}\big[\ln q(\bar{\mathbf{x}})\big]-\mathbb{E}_{q(\bar{\mathbf{x}})}\big[\ln p(\bar{\mathbf{x}}|\bar{\mathbf{y}})\big]\big\},$
where $q^{\star}(\bar{\mathbf{x}})$ denotes the optimal variational distribution. To approximate the posterior, we adopt the mean-field VB inference method, for which the variational distribution $q(\bar{\mathbf{x}})$ is drawn from a mean-field variational family [53], such that
$q(\bar{\mathbf{x}})=\prod_{i=1}^{V}q_{i}(x_{i}),$ (23)
where the latent variables are mutually independent, with each being governed by a distinct factor in the variational density. The optimal value of $q_{i}(x_{i})$ is obtained as [53, Chapter 10]
$q_{i}^{\star}(x_{i})$ $\propto\mathrm{exp}\left\{\big\langle{\ln p(\bar{\mathbf{y}},\bar{\mathbf{x}})\big\rangle}\right\},$ (24)
where $\langle\cdot\rangle$ is computed utilizing the currently fixed variational density $q_{-i}(\bar{\mathbf{x}}_{-i})=\prod_{j=1,j\neq i}^{V}q_{j}(x_{j})$. To optimize (22), we use the Coordinate Ascent Variational Inference algorithm, which is an iterative method that ensures convergence to at least a locally optimal solution [54].
III-B VB Inference for Few-bit $\Sigma\Delta$ MIMO Detection
@@
This integral is intractable to evaluate in closed-form due to the high-dimensional integration and the intricate correlation between $\mathbf{r}$ and $\mathbf{s}$. This challenge motivates us to explore a novel VB inference method to solve the data detection problem with $\Sigma\Delta$ arrays.
@@
We consider the residual interference-plus-noise as an unknown parameter $N_{0}^{\rm post}$, which must be estimated using the VB method. The dependency between random variables under spatial $\Sigma\Delta$ processing can be graphically modeled through the Bayesian network in Fig. 3, where $\gamma\triangleq 1/N_{0}^{\rm post}$ denotes the precision that is floated as an unknown random variable and the arrows indicate the conditional probability between variables. The main goal is to infer the distribution of data $\mathbf{s}$ given the observation $\mathbf{y}$. To accomplish this, we employ the mean-field VB inference method, which derives the mean-field variational distributions $q(r_{i})$, $q(s_{k})$, and $q(\gamma)$, from the factorized family to approximate the posterior for individual variables, such that
@@
$\;\;\times\prod_{i=1}^{N}p(y_{i}|r_{i})p(r_{i}|r_{i-1},y_{i-1},\mathbf{s},\gamma;\mathbf{H}),$ (27)
@@
1) Updating $r_{i}$: For $i=1,\ldots,N-1$, the variational distribution $q(r_{i})$ can be obtained by taking the expectation of the conditional in (III-B) w.r.t. $q(\mathbf{s},\gamma)$ as
@@
Defining $v_{i}\triangleq(a_{i}+b_{i})/2$, we see that the variational distribution $q(r_{i})$ in (III-B) is a truncated complex Gaussian distribution obtained by pruning the distribution $\mathcal{CN}\big(v_{i},1/(2\langle\gamma\rangle)\big)$ onto the interval $y_{i}^{\mathrm{low}},y_{i}^{\mathrm{up}}]$. In Appendix B we show that the mean and variance of the resulting distribution are given by $\langle r_{i}\rangle=\mathsf{F}_{r}(v_{i},2\langle\gamma\rangle,y_{i}^{\mathrm{low}},y_{i}^{\mathrm{up}})$} and $\tau_{r_{i}}={\sf G}_{r}(v_{i},2\langle\gamma\rangle,y_{i}^{\mathrm{low}},y_{i}^{\mathrm{up}})$}, respectively.
For $i=N$, while the factor $p(r_{N+1}|r_{N},y_{N},\mathbf{s},\gamma;\mathbf{H})$ does not exist, the variational distribution $q(r_{N})$ can be computed similarly to (III-B) and is given by
@@
2) Updating $s_{k}$: The variational distribution $q(s_{k})$ is obtained by taking the expectation of the conditional in (III-B) w.r.t. $q(\mathbf{r},\gamma)$:
@@
$\mathbf{z}_{k}=\mathbf{h}_{k}s_{k}+\tilde{\mathbf{n}}_{k},$ (34)
where $\tilde{\mathbf{n}}_{k}\sim\mathcal{CN}(\mathbf{0},\langle\gamma\rangle^{-1}\mathbf{I}_{N})$ and $\mathbf{z}_{k}=[z_{1,k},\ldots,z_{N,k}]^{T}$. The variational mean and variance of $s_{k}$ can be expressed as $\langle s_{k}\rangle={\sf F}_{s}(\mathbf{z}_{k},\mathbf{h}_{k},\langle\gamma\rangle)$ and $\tau_{s_{k}}={\sf G}_{s}(\mathbf{z}_{k},\mathbf{h}_{k},\langle\gamma\rangle)$, respectively, as shown in Appendix C.
3) Updating $\gamma$: We assume a Gamma distribution $p(\gamma)={\rm Gamma}(\alpha,\beta)$, as a conjugate prior for the precision parameter $\gamma$, where $\alpha$ and $\beta$ are the shape and rate parameters, respectively. This enables the derivation of the variational distribution $q(\gamma)$ in closed form. The choice of a Gamma distribution for $\gamma$ is motivated by its role as a conjugate prior for the precision of a Gaussian likelihood in the noise model (1), and this conjugate prior is proportional to the product of a power of $\gamma$ and the exponential of a linear function of $\gamma$ [53, Chapter 2, Section 2.3.6]. Taking the expectation of the conditional distribution in (III-B) w.r.t. $q(\mathbf{s},\mathbf{r})$, $q(\gamma)$ can be obtained as
@@
[…]
Given that we have assigned a conjugate Gamma prior $p(\gamma)={\rm Gamma}(\alpha,\beta)$, the resulting variational distribution follows a Gamma distribution. Specifically, after computing the expectation and simplifying the exponent, $q(\gamma)$ is a Gamma distribution with mean
$\langle\gamma\rangle=\frac{N+\alpha}{\beta+\|\mathbf{u}\|^{2}+2\mathrm{Tr}\{\bm{\Sigma}_{\mathbf{r}}\}-\tau_{r_{N}}+\mathrm{Tr}\big\{\mathbf{H}\bm{\Sigma}_{\mathbf{s}}\mathbf{H}^{H}\big\}},$ (37)
where $N+\alpha$ reflects the updated shape parameter, and the denominator incorporates the residual noise $\|\mathbf{u}\|^{2}$, the trace of the variance of $\mathbf{r}$, and the channel-data covariance term, consistent with the noise model and dependencies in the $1^{\rm st}$-order $\Sigma\Delta$ architecture.
By iteratively optimizing $q(r_{i})$, $q(s_{k})$, and $q(\gamma)$ through the updates of $\langle x_{i}\rangle$}, $\langle s_{k}\rangle$}, and $\langle\gamma\rangle$, we derive the CAVI algorithm, as presented in Algorithm 1, for the $1^{\rm st}$-order $\Sigma\Delta$ quantization. We designate this algorithm as the $1^{\rm st}$-order SD-VB algorithm, where the parameter $\epsilon$ is the numerator of $\hat{\gamma}$ in the variational distributions of (III-B) and (30) for updating $r_{i}$ and $r_{N}$, respectively.
Remark 1: In Algorithm 1, we denote the estimates of the variational means $\langle r_{i}\rangle,\langle s_{k}\rangle,\langle\gamma\rangle$ at iteration $t$ as $\hat{r}_{i}^{t},\hat{s}_{k}^{t},\hat{\gamma}^{t}$, respectively, and each iteration includes a round of updating the estimates of $\mathbf{r},\mathbf{s}$, and $\gamma$. These estimates, computed via closed-form updates, represent the mean values of $q(r_{i})$, $q(s_{k})$, and $q(\gamma)$, driving the algorithm toward convergence. Recalling $u_{i}=\langle r_{i}\rangle-\mathrm{e}^{-j\phi}(\langle r_{i-1}\rangle-y_{i-1})-\mathbf{h}_{i}^{T}\langle\mathbf{s}\rangle$, $a_{i}=\mathbf{h}_{i}^{T}\langle\mathbf{s}\rangle+\mathrm{e}^{-j\phi}(\langle r_{i-1}\rangle-y_{i-1})$ and $b_{i}=y_{i}+\mathrm{e}^{j\phi}\langle r_{i+1}\rangle-\mathrm{e}^{j\phi}\mathbf{h}_{i+1}^{T}\langle\mathbf{s}\rangle$, $a_{i}$ and $b_{i}$ can be rewritten as $a_{i}=\hat{r}_{i}^{t}-u_{i}$ and $b_{i}=\hat{r}_{i}^{t}+\mathrm{e}^{j\phi}u_{i+1}$ using the current estimates $\big\{\hat{r}_{i}^{t}\big\}$ and $\big\{\hat{\mathbf{s}}^{t}\big\}$, where $\{\hat{r}_{i}^{t}\}$ denotes the set of estimated pre-quantized signals $\hat{r}_{i}^{t}$, across all antennas (i=1,,Ni=1,...,N) at iteration $t$. Thus, we have
@@
IV Extension to MIMO Detection with $2^{\rm nd}$-Order $\Sigma\Delta$ Quantization